# Mixed-state geometric phases of coherent and squeezed spin states

### v1

@@ Abstract
Two mixed-state geometric phases, known as the Uhlmann phase and interferometric phase (IGP), of spin coherent states (CSSs) and spin squeezed states (SSSs) are analyzed. While each phase follows its parallel-transport condition, we also consider the non-adiabatic IGP for arbitrary unitary evolution beyond parallel transport. For the $j=3/2$ CSS, the Uhlmann phase shows temperature-induced topological phase transitions with jumps. The IGP and non-adiabatic IGP for the $j=3/2$ CSS also exhibits temperature-induced jumps. In contrast, the Uhlmann phase of the $j=1$ SSS exhibits smooth behavior without any temperature-induced transition. Interestingly, the parallel-transport condition of the IGP of the $j=1$ SSS in general does not allow a solution at finite temperature. Instead, the non-adiabatic IGP for the $j=1$ SSS has a solution showing smooth behavior as the squeezing parameter and temperature change. We also briefly discuss possible experimental implications and simulations.
@@ I Introduction
Geometric phases have been a key concept in topological properties of quantum systems [1, 2]. The Berry phase arises when a quantum system undergoes adiabatic cyclic evolution in the context of pure quantum states and has been fundamental in studying topological insulators and superconductors, where the Berry curvature characterizes the underlying topological features [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. However, quantum systems are often in mixed states due to finite temperatures or non-equilibrium conditions, which call for generalizations of the Berry phase to mixed quantum states. A mathematically rigorous approach is the Uhlmann phase [17, 18, 19, 20] for mixed states. Like the Berry phase, the Uhlmann phase is obtained from parallel transport in a cyclic process, and the Uhlmann-Berry correspondence shows that the Uhlmann phase in the low-temperature limit generally agrees with the Berry phase of the ground state [21]. However, the Uhlmann bundle is trivial [22], which limits its applicability to defining other quantized topological indices. Nevertheless, the Uhlmann phase of bosonic and fermionic coherent states have been studied [21], in addition to that of two-level and spin-$j$ systems [23, 24, 25]. With a suitable generalization, the Uhlmann phase of dynamic systems has also been studied [26].
@@
Another approach called the interferometric geometric phase (IGP) for mixed states has been proposed [27] and further analyzed [28, 29, 30, 31, 32], extending the geometric phase concept via an analogy of the optical Mach-Zehnder interferometer [33]. This IGP has been experimentally observed using nuclear magnetic resonance and polarized neutrons [34, 35, 36, 33]. Unlike the Uhlmann phase, the IGP does not require the evolution of the system to be cyclic [27]. In some cases, it can be shown that the IGP is the argument of the weighted sum of the Berry-phase factors of the constituent states [27].
@@
Ref. [37] presents a comparison between the Uhlmann phase and IGP to contrast their differences. The Uhlmann phase, for instance, requires the ancillary system modeling environmental effects to be manipulated according to the change of the system and can exhibit finite-temperature topological phase transitions with quantized jumps in two-level systems. In contrast, the IGP does not exhibit such transitions at finite temperatures for general two-level systems. Recent studies have demonstrated that the IGP remains intact as temperature varies for the Kitaev chain [38], in contrast to the discrete jumps seen in the Uhlmann phase [23]. Nevertheless, quantized jumps of the IGP as temperature increases can occur in three-level or higher-dimensional systems [37]. Interestingly, these generalizations of the mixed-state geometric phases provide further insights into topological quantum phase transitions (TQPTs) and dynamical quantum phase transitions (DQPTs) [39].
@@
Meanwhile, spin systems have served as paradigms in various physics branches, ranging from quantum mechanics [40] to statistical physics [41] to spintronic devices [42]. Two types of spin states have received particular attention: The coherent spin states (CSSs) are minimal uncertainty states of collective angular momentum. The CSS is the building block for the path-integral description of many-body spin systems [43] and serves as a classical approximation of quantum spin systems, thereby establishing the standard quantum limit (SQL) for measurements [44, 45, 46, 47]. However, surpassing the SQL is critical for advancing quantum metrology and quantum information technologies. Spin squeezing [48, 49, 50] allows for a reduction of the variance of one spin component at the expense of increasing that of another orthogonal direction, resulting in the squeezed spin states (SSSs). The SSSs enable precision measurements beyond the SQL and approach the Heisenberg limit with potential applications in, for example, atomic clocks [51, 52]. Significant improvements in precision measurement using spin squeezing have also been made in atomic Bose-Einstein condensates (BECs) [53, 54]. These developments not only highlight the metrological advantage of the SSS but also open possible avenues for generating quantum entanglement in many-body systems, fostering further advancement in quantum information processing [55].
@@
Complementing the progress in surpassing the SQL through spin squeezing, here we set to investigate the geometric phases of the CSS and SSS at finite temperatures and search for topological phases or transitions. The possibility of finding different mixed-state geometric phases according to different evolution conditions already demonstrates rich physics in the underlying topology of finite-temperature systems. While the geometric phase of the ground-state CSS has been analyzed [56, 57], here we explicitly calculate the Uhlmann phase and the IGP for selected CSS and SSS at finite temperatures. We found that for the $j=3/2$ CSS, both Uhlmann phase and IGP demonstrate temperature-dependent topological or geometric phase transitions with quantized jumps. The jumps survive after generalizing the IGP to the non-adiabatic IGP for the CSS. In contrast, the $j=1$ SSS has a smoothly-varying Uhlmann phase as temperature increases, so there is no topological phase transition. More interestingly, the parallel-transport condition of the $j=1$ SSS cannot be satisfied at finite temperatures in general, leaving a construction of the IGP impractical. We therefore analyze the non-adiabatic IGP for the $j=1$ SSS to circumvent the obstacle but found no discrete jump in the geometric contribution of the phase at finite temperature.
@@ II.1 Overview of CSS
The spin operators satisfy the algebra [40]
@@
The eigenstates of $J^{2}$ and $J_{z}$ are $|jm\rangle,\leavevmode\nobreak\ -j\leq m\leq j$. Moreover, $J_{\pm}|jm\rangle=\sqrt{(j\mp m)(j\pm m+1)}|jm\pm 1\rangle$ and $J_{z}|jm\rangle=m|jm\rangle$. The CSS is defined as [58, 59]
@@
Using the Baker-Campbell-Hausdorff disentangling formula [58], the CSS can also be expressed as
@@
with $\zeta=\zeta(\xi)=\frac{\xi\tan|\xi|}{|\xi|}$. After defining $D(\xi)\equiv D(\zeta)=\mathrm{e}^{\xi J_{+}-\bar{\xi}J_{-}}$, the CSS corresponds to the ground state of the translated Hamiltonian $\hat{H}(\xi)\equiv\hat{H}(\zeta)=D(\zeta)\hat{H}D^{\dagger}(\zeta)$. Here $\hat{H}=\omega_{0}J_{z}$ , and $\omega_{0}$ is the Larmor frequency. The associated excited states are obtained in a similar manner: $|jm,\xi\rangle=|jm,\zeta\rangle=D(\zeta)|jm\rangle$, $-j\leq m\leq j$. By parameterizing $\xi$ as $\xi=\mathrm{e}^{-\mathrm{i}\phi}\frac{\theta}{2}$, we have $\zeta=\mathrm{e}^{-\mathrm{i}\phi}\tan\frac{\theta}{2}$. If the parameter ranges are chosen as $0\leq\phi\leq 2\pi$, $0\leq\theta\leq\pi$, then $\zeta$ spans the entire complex plane. Consequently, we will use $\zeta$ as the control parameter. The map $(\theta,\phi)\rightarrow\zeta$ projects the unit sphere onto the entire complex plane. Actually, the CSSs are closely related to the eigenstates of the corresponding spin-$j$ system with the Hamiltonian $H=\omega_{0}\hat{\mathbf{B}}\cdot\mathbf{J}$, where $\hat{\mathbf{B}}$ is the unit vector pointing along the direction of an external magnetic field [43].
@@ II.2 Overview of SSS
where $\sigma_{l\alpha}$ is the Pauli matrix for the $l$-th object. Here we consider the spin-squeezed states (SSS) generated by the one-axis twisting Hamiltonian [49, 60]
@@ II.3 Uhlmann phase
A mathematically rigorous approach for generalizing the Berry phase from pure states to mixed states is the Uhlmann phase [17] through purification of density matrices, which can be described by the underlying fiber-bundle structure [61], An operator $W$ is said to purify the density matrix $\rho$ if $\rho=WW^{\dagger}$. Conversely, the purification $W$ can be uniquely decomposed as $W=\sqrt{\rho}\mathcal{U}$ if $\rho$ is full-rank (i.e., rank$(\rho)=N$ for a system with a $N$-dimensional Hilbert space). Here the unitary operator $\mathcal{U}$ is referred to as the phase factor of mixed states, representing a generalization of the U(1) phase factor associated with pure states.
We consider a system depending on a set of parameters $\mathbf{R}=(R_{1},R_{2},\cdots,R_{k})^{T}\in M$, where $M$ is the parameter manifold. When $\rho$ evolves along a smooth curve $C(t):=\mathbf{R}(t)$ in $M$, the purification of the density matrix also continuously evolves along an induced curve $\gamma(t):=W(t)=\sqrt{\rho(\mathbf{R}(t))}\mathcal{U}(\mathbf{R}(t))$. If the length of $\gamma$, defined as $L(\gamma)={\int}_{\gamma}\sqrt{\text{Tr}(\dot{W}\dot{W}^{\dagger})}\mathrm{d}t$, is minimized, $W$ is said to undergo a parallel-transport process. Therefore, $\gamma(t)$ is the horizontal lift of $\rho(t)\equiv\rho(\mathbf{R}(t))$. It can be shown that the Uhlmann parallel-transport condition is given by [62]
@@
is the Uhlmann connection. Here $\lambda_{n}$, $|n\rangle$ represent the eigenvalues and eigenstates of $\rho$, respectively. Notably, the parallelity lacks transitivity [39] even during parallel transport. Hence, the operator $\mathcal{P}\mathrm{e}^{-\oint_{C}A_{U}}$ associated with the Uhlmann holonomy quantifies a measure of the loss of parallelity. The following Uhlmann phase then reflects the topological change of the Uhlmann holonomy.
@@ II.4 Interferometric geometric phase
There is another approach for constructing a geometric phase for mixed states. Inspired by optical processes in the Mach-Zehnder interferometer, Ref. [27] assigned a phase to a mixed quantum state represented by the density matrix $\rho$ undergoing a unitary evolution $U(t)$. Explicitly, if a system evolves according to $\rho(t)=U(t)\rho(0)U^{\dagger}(t)$, it gives rise to a total phase
@@
Notably, the transformation from $\rho(t)$ to $\rho(t+\mathrm{d}t)$ is $\rho(t+\mathrm{d}t)=U(t+\mathrm{d}t)\rho(0)U^{\dagger}(t+\mathrm{d}t)=U(t+\mathrm{d}t)U^{\dagger}(t)\rho(t)U(t)U^{\dagger}(t+\mathrm{d}t)$. Thus, by definition, the relative phase between the mixed states at $t$ and $t+\mathrm{d}t$ is $\theta_{\text{T}}(t)=\arg\text{Tr}\left[\rho(t)U(t+\mathrm{d}t)U^{\dagger}(t)\right]$. In this framework [27], a parallel-transport condition of $U(t)$ is proposed as follows. $\rho(t+\mathrm{d}t)$ is required to be instantaneously “in phase” with $\rho(t)$, which means $\arg\text{Tr}\left[\rho(t)U(t+\mathrm{d}t)U^{\dagger}(t)\right]=0$. In the differential form, the condition yields
@@
To specify parallel-transport by the evolution according to $U(t)$ more accurately, the condition (15) has been strengthened as [27]
@@
where $|n(t)\rangle$ is the $n$th eigen-vector of $\rho(t)$. We emphasize that the above parallel-transport condition is distinct from the Uhlmann parallel-transport condition (10) which depends explicitly on the choice of the U$(N)$ phase factor $\mathcal{U}$.
@@
As a consequence, the total phase contains only the geometric contribution from the IGP.
In certain situations, it may be difficult for unitary evolutions to fully satisfy the parallel-transport condition (15). Nevertheless, the above discussion can be extended to general unitary evolution beyond parallel transport, which results in non-adiabatic phases [63]. For pure states, a similar formalism is known as the quantum kinematic approach to the geometric phase [64]. Recently, a similar non-adiabatic geometric phase for mixed states has been discussed in Ref. [65]. We follow those ideas and define the non-adiabatic IGP for any unitary process by subtracting the non-adiabatic generalization of the dynamic phase from the total phase. Explicitly,
@@
II.5 Difference between the two geometrical phases
We emphasize that the two aforementioned geometric phases for mixed states are physically and mathematically distinct. Although the theory of IGP can also be formulated equivalently using a purification formalism of density matrices [27], any phase factor $\mathcal{U}$ is ruled out in the derivation. Specifically, only the trivial decomposition $W=\sqrt{\rho}$ is permissible for the IGP because otherwise the condition (15) would be violated.
In contrast, the generation of the Uhlmann phase inherently relies on the phase factor $\mathcal{U}$ according to the underlying fiber-bundle structure. All purifications of a given density matrix $\rho$ thus span a U$(N)$ fiber space. Consequently, the Uhlmann phase is naturally and directly connected to the topological properties of the U$(N)$ principal bundle, a feature absent in the IGP. Explicitly, the Uhlmann phase corresponds to the Uhlmann holonomy of the Uhlmann bundle of density matrices, similar to the Berry phase revealing the Berry holonomy of the fiber bundle of pure states [61]. Moreover, the Uhlmann phase requires a cyclic process but is guaranteed to be gauge invariant [27, 37] while the IGP and its generalization do not impose those constraints. In the following, we will compare the differences between the Uhlmann phase and IGP through specific examples of the CSS and SSS.

### v2

@@ Abstract
Two mixed-state geometric phases, known as the Uhlmann phase and interferometric phase (IGP), of spin coherent states (CSSs) and spin squeezed states (SSSs) are analyzed. While each phase follows its parallel-transport condition, we also consider the non-adiabatic IGP for arbitrary unitary evolution beyond parallel transport. For the $j=3/2$ CSS, the Uhlmann phase shows temperature-induced topological phase transitions with jumps. The IGP and non-adiabatic IGP for the $j=3/2$ CSS also exhibits temperature-induced jumps. In contrast, the Uhlmann phase of the $j=1$ one-axis SSS exhibits smooth behavior without any temperature-induced transition. Interestingly, the parallel-transport condition of the IGP of the $j=1$ one-axis SSS only admits a trivial solution at finite temperature while the non-adiabatic IGP for the $j=1$ one-axis SSS has a solution showing smooth behavior as the squeezing parameter and temperature change. By considering $j=1$ two-axis SSS, the IGP and its non-adiabatic generalization exhibit smooth behavior at finite temperatures. We also briefly discuss possible experimental implications and simulations.
@@ I Introduction
Geometric phases have been a key concept in topological properties of quantum systems Bohm et al. (2003a); Chruscinski and Jamiolkowski (2004). The Berry phase arises when a quantum system undergoes adiabatic cyclic evolution in the context of pure quantum states and has been fundamental in studying topological insulators and superconductors, where the Berry curvature characterizes the underlying topological features Thouless et al. (1982); Haldane (1988); Hasan and Kane (2010); Qi and Zhang (2011); Kane and Mele (2005a, b); Chiu et al. (2016); Bernevig and Hughes (2013); Bernevig and Zhang (2006); Moore and Balents (2007); Fu et al. (2007); Bohm et al. (2003b); Vanderbilt (2018); Cohen et al. (2019). However, quantum systems are often in mixed states due to finite temperatures or non-equilibrium conditions, which call for generalizations of the Berry phase to mixed quantum states. A mathematically rigorous approach is the Uhlmann phase Uhlmann (1986, 1989, 1992, 1991) for mixed states. Similar to the Berry phase associated with the holonomy of the Berry bundle, the Uhlmann phase is obtained from parallel transport in a cyclic process and reflects the holonomy of the Uhlmann bundle. Moreover, the Uhlmann-Berry correspondence shows that the Uhlmann phase in the low-temperature limit generally agrees with the Berry phase of the ground state Wang et al. (2023). However, the Uhlmann bundle is trivial Budich and Diehl (2015), which limits its applicability to defining other quantized topological indices. Nevertheless, the Uhlmann phase of bosonic and fermionic coherent states have been studied Wang et al. (2023), in addition to that of two-level and spin-$j$ systems Viyuela et al. (2014); Hou et al. (2021a); Morachis Galindo et al. (2021). With a suitable generalization, the Uhlmann phase of dynamic systems has also been studied Tang et al. (2024).
@@
Another approach called the interferometric geometric phase (IGP) for mixed states has been proposed Sjöqvist et al. (2000) and further analyzed Ericsson et al. (2003); Sjöqvist (2004); de Faria et al. (2003); Chaturvedi et al. (2004); Tong et al. (2004); Kwek et al. (2006), extending the geometric phase concept via an analogy of the optical Mach-Zehnder interferometer Ericsson et al. (2005). This IGP has been experimentally observed using nuclear magnetic resonance and polarized neutrons Du et al. (2003); Ghosh and Kumar (2006); Klepp et al. (2008); Ericsson et al. (2005). Unlike the Uhlmann phase, the IGP does not require the evolution of the system to be cyclic Sjöqvist et al. (2000). In some cases, it can be shown that the IGP is the argument of the weighted sum of the Berry-phase factors of the constituent states Sjöqvist et al. (2000).
@@
Ref. Hou et al. (2023a) presents a comparison between the Uhlmann phase and IGP to contrast their differences. The Uhlmann phase, for instance, requires the ancillary system modeling environmental effects to be manipulated according to the change of the system and can exhibit finite-temperature topological phase transitions with quantized jumps in two-level systems. In contrast, the IGP does not exhibit such transitions at finite temperatures for general two-level systems. Recent studies have demonstrated that the IGP remains intact as temperature varies for the Kitaev chain Andersson et al. (2016), in contrast to the discrete jumps seen in the Uhlmann phase Viyuela et al. (2014). Nevertheless, quantized jumps of the IGP as temperature increases can occur in three-level or higher-dimensional systems Hou et al. (2023a). Interestingly, these generalizations of the mixed-state geometric phases provide further insights into topological quantum phase transitions (TQPTs) and dynamical quantum phase transitions (DQPTs) Hou et al. (2020).
@@
Meanwhile, spin systems have served as paradigms in various physics branches, ranging from quantum mechanics Sakurai and Napolitano (2010) to statistical physics Huang (1991) to spintronic devices Bandyopadhyay and Cahay (2015). Two types of spin states have received particular attention: The coherent spin states (CSSs) are minimal uncertainty states of collective angular momentum. The CSS is the building block for the path-integral description of many-body spin systems Auerbach (1994) and serves as a classical approximation of quantum spin systems, thereby establishing the standard quantum limit (SQL) for measurements Caves (1981); Giovannetti et al. (2004); Wineland et al. (1992); Pezze and Smerzi (2014). However, surpassing the SQL is critical for advancing quantum metrology and quantum information technologies. Spin squeezing Kitagawa and Ueda (1993); Ma et al. (2011); Block et al. (2024) allows for a reduction of the variance of one spin component at the expense of increasing that of another orthogonal direction, resulting in the squeezed spin states (SSSs). The SSSs enable precision measurements beyond the SQL and approach the Heisenberg limit with potential applications in, for example, atomic clocks Wineland et al. (1994); Eckner et al. (2023). Significant improvements in precision measurement using spin squeezing have also been made in atomic Bose-Einstein condensates (BECs) Estève et al. (2008); Gross et al. (2010). These developments not only highlight the metrological advantage of the SSS but also open possible avenues for generating quantum entanglement in many-body systems, fostering further advancement in quantum information processing Riedel et al. (2010).
@@
Complementing the progress in surpassing the SQL through spin squeezing, here we set to investigate the geometric phases of the CSS and SSS at finite temperatures and search for topological phases or transitions. The possibility of finding different mixed-state geometric phases according to different evolution conditions already demonstrates rich physics in the underlying topology of finite-temperature systems. While the geometric phase of the ground-state CSS has been analyzed Chaturvedi et al. (1987); Chryssomalakos et al. (2018), here we explicitly calculate the Uhlmann phase and the IGP for selected CSS and SSS at finite temperatures. We found that for the $j=3/2$ CSS, both Uhlmann phase and IGP demonstrate temperature-dependent topological or geometric phase transitions with quantized jumps. The jumps survive after generalizing the IGP to the non-adiabatic IGP for the CSS. In contrast, the $j=1$ SSS has a smoothly-varying Uhlmann phase as temperature increases, so there is no topological phase transition. Moreover, the parallel-transport condition of the one-axis $j=1$ SSS leads to a trivial solution finite temperatures in general, leaving no feature of the IGP. By considering two-axis squeezing, the IGP and non-adiabatic IGP of the resulting SSS exhibit smooth behavior at finite temperatures. We therefore analyze the non-adiabatic IGP for the $j=1$ SSS to circumvent the obstacle but found no discrete jump in the geometric contribution of the phase at finite temperature.
@@ II.1 Overview of CSS
The spin operators satisfy the algebra Sakurai and Napolitano (2010)
@@
The eigenstates of $J^{2}$ and $J_{z}$ are $|jm\rangle,\leavevmode\nobreak\ -j\leq m\leq j$. Moreover, $J_{\pm}|jm\rangle=\sqrt{(j\mp m)(j\pm m+1)}|jm\pm 1\rangle$ and $J_{z}|jm\rangle=m|jm\rangle$. The CSS is defined as Perelomov (1986); Radcliffe (1971)
@@
Using the Baker-Campbell-Hausdorff disentangling formula Perelomov (1986), the CSS can also be expressed as
@@
with $\zeta=\zeta(\xi)=\frac{\xi\tan|\xi|}{|\xi|}$. After defining $D(\xi)\equiv D(\zeta)=\mathrm{e}^{\xi J_{+}-\bar{\xi}J_{-}}$, the CSS corresponds to the ground state of the translated Hamiltonian $\hat{H}(\xi)\equiv\hat{H}(\zeta)=D(\zeta)\hat{H}D^{\dagger}(\zeta)$. Here $\hat{H}=\omega_{0}J_{z}$ , and $\omega_{0}$ is the Larmor frequency. The associated excited states are obtained in a similar manner: $|jm,\xi\rangle=|jm,\zeta\rangle=D(\zeta)|jm\rangle$, $-j\leq m\leq j$. By parameterizing $\xi$ as $\xi=\mathrm{e}^{-\mathrm{i}\phi}\frac{\theta}{2}$, we have $\zeta=\mathrm{e}^{-\mathrm{i}\phi}\tan\frac{\theta}{2}$. If the parameter ranges are chosen as $0\leq\phi\leq 2\pi$, $0\leq\theta\leq\pi$, then $\zeta$ spans the entire complex plane. Consequently, we will use $\zeta$ as the control parameter. The map $(\theta,\phi)\rightarrow\zeta$ projects the unit sphere onto the entire complex plane. Actually, the CSSs are closely related to the eigenstates of the corresponding spin-$j$ system with the Hamiltonian $H=\omega_{0}\hat{\mathbf{B}}\cdot\mathbf{J}$, where $\hat{\mathbf{B}}$ is the unit vector pointing along the direction of an external magnetic field Auerbach (1994).
@@ II.2 Overview of SSS
where $\sigma_{l\alpha}$ is the Pauli matrix for the $l$-th object. Here we consider the spin-squeezed states (SSS) generated by the one-axis twisting Hamiltonian Ma et al. (2011); Xue et al. (2007)
@@
Another spin squeezing method is based on a generalized two-axis counter-twisting Hamiltonian Kitagawa and Ueda (1993); Ma et al. (2011), which reads
$H_{\mathrm{TAT}}=\mathrm{i}(zJ^{2}_{+}-\bar{z}J^{2}_{-}).$ (10)
Here $z\in\mathds{C}$ is a complex number which can be parameterized as $z=\mathrm{e}^{-\mathrm{i}\phi}\tan\left(\frac{\theta}{2}\right)$ with the parameter ranges chosen as $0\leq\phi\leq 2\pi$, $0\leq\theta\leq\pi$. This is a quadratic squeezing Hamiltonian and it will degrade to the ordinary two-axis Hamiltonian $\frac{\gamma}{2\mathrm{i}}(J_{+}^{2}-J_{-}^{2})$ if $z$ is a purely real number equal to $-\gamma/2$. The squeezing operator $K(z)$ now is defined as
$K(z)=\mathrm{e}^{zJ^{2}_{+}-\bar{z}J^{2}_{-}},$ (11)
which is somewhat similar with the translation operator $D(\xi)=\mathrm{e}^{\xi J_{+}-\bar{\xi}J_{-}}$ we encountered in the discussion of the CSS. However, there are subtle differences in the quadratic form of angular momentum operators. Similar to the case of one-axis squeezing, the eigenstates $|jm,z\rangle$ can be given from the corresponding Hamiltonian
$H(z)=K(z)\omega_{0}J_{z}K^{\dagger}(z).$ (12)
@@ II.3 Uhlmann phase
A mathematically rigorous approach for generalizing the Berry phase from pure states to mixed states is the Uhlmann phase Uhlmann (1986) through purification of density matrices, which can be described by the underlying fiber-bundle structure Guo et al. (2020), An operator $W$ is said to purify the density matrix $\rho$ if $\rho=WW^{\dagger}$. Conversely, the purification $W$ can be uniquely decomposed as $W=\sqrt{\rho}\mathcal{U}$ if $\rho$ is full-rank (i.e., rank$(\rho)=N$ for a system with a $N$-dimensional Hilbert space). Here the unitary operator $\mathcal{U}$ is referred to as the phase factor of mixed states, representing a generalization of the U(1) phase factor associated with pure states.
We consider a system depending on a set of parameters $\mathbf{R}=(R_{1},R_{2},\cdots,R_{k})^{T}\in M$, where $M$ is the parameter manifold. When $\rho$ evolves along a smooth curve $C(t):=\mathbf{R}(t)$ in $M$, the purification of the density matrix also continuously evolves along an induced curve $\gamma(t):=W(t)=\sqrt{\rho(\mathbf{R}(t))}\mathcal{U}(\mathbf{R}(t))$. If the length of $\gamma$, defined as $L(\gamma)={\int}_{\gamma}\sqrt{\text{Tr}(\dot{W}\dot{W}^{\dagger})}\mathrm{d}t$, is minimized, $W$ is said to undergo a parallel-transport process. Therefore, $\gamma(t)$ is the horizontal lift of $\rho(t)\equiv\rho(\mathbf{R}(t))$. It can be shown that the Uhlmann parallel-transport condition is given by Uhlmann (1995)
@@
is the Uhlmann connection. Here $\lambda_{n}$, $|n\rangle$ represent the eigenvalues and eigenstates of $\rho$, respectively. Notably, the parallelity lacks transitivity Hou et al. (2020) even during parallel transport. Hence, the operator $\mathcal{P}\mathrm{e}^{-\oint_{C}A_{U}}$ associated with the holonomy of the Uhlmann bundle quantifies a measure of the loss of parallelity between the initial and final purifications. The following Uhlmann phase then reflects the topological change of the Uhlmann holonomy.
@@ II.4 Interferometric geometric phase
There is another approach for constructing a geometric phase for mixed states. Inspired by optical processes in the Mach-Zehnder interferometer, Ref. Sjöqvist et al. (2000) assigned a phase to a mixed quantum state represented by the density matrix $\rho$ undergoing a unitary evolution $U(t)$. Explicitly, if a system evolves according to $\rho(t)=U(t)\rho(0)U^{\dagger}(t)$, it gives rise to a total phase
@@
Notably, the transformation from $\rho(t)$ to $\rho(t+\mathrm{d}t)$ is $\rho(t+\mathrm{d}t)=U(t+\mathrm{d}t)\rho(0)U^{\dagger}(t+\mathrm{d}t)=U(t+\mathrm{d}t)U^{\dagger}(t)\rho(t)U(t)U^{\dagger}(t+\mathrm{d}t)$. Thus, by definition, the relative phase between the mixed states at $t$ and $t+\mathrm{d}t$ is $\theta_{\text{T}}(t)=\arg\text{Tr}\left[\rho(t)U(t+\mathrm{d}t)U^{\dagger}(t)\right]$. In this framework Sjöqvist et al. (2000), a parallel-transport condition of $U(t)$ is proposed as follows. $\rho(t+\mathrm{d}t)$ is required to be instantaneously “in phase” with $\rho(t)$, which means $\arg\text{Tr}\left[\rho(t)U(t+\mathrm{d}t)U^{\dagger}(t)\right]=0$. In the differential form, the condition yields
@@
To specify parallel-transport by the evolution according to $U(t)$ more deterministically, the condition (18) has been strengthened as Sjöqvist et al. (2000)
@@
where $|n(t)\rangle$ is the $n$th eigen-vector of $\rho(t)$. We emphasize that the parallel-transport condition (18) for the IGP fundamentally differs from Uhlmann’s condition in Eq. (13), due to the U$(N)$ gauge dependence of the phase factor $\mathcal{U}$. Crucially, introducing $\mathcal{U}$ via $W=\sqrt{\rho}\mathcal{U}$ can violate the parallel-transport condition (18) Hou et al. (2023b).
@@
As a consequence, the total phase contains only the geometric contribution from the IGP. We note that the strengthened condition (19) will result in vanishing integrand in the above expression, thereby resulting the same vanishing dynamic phase.
In certain situations, it may be difficult for unitary evolutions to fully satisfy the parallel-transport condition (18). Nevertheless, the above discussion can be extended to general unitary evolution beyond parallel transport, which results in non-adiabatic phases Zhang et al. (2023). For pure states, a similar formalism is known as the quantum kinematic approach to the geometric phase Mukunda and Simon (1993). Recently, a similar non-adiabatic geometric phase for mixed states has been discussed in Ref. Mohammadi Almas and Najarbashi (2024). We follow those ideas and define the non-adiabatic IGP for any unitary process by subtracting the non-adiabatic generalization of the dynamic phase from the total phase. Explicitly,
@@
II.5 Difference between the two geometric phases
We emphasize that the two aforementioned geometric phases for mixed states are physically and mathematically distinct. Although the theory of IGP can also be formulated equivalently using a purification formalism of density matrices Sjöqvist et al. (2000), any phase factor $\mathcal{U}$ is ruled out in the derivation. Specifically, only the trivial decomposition $W=\sqrt{\rho}$ is permissible for the IGP because otherwise the condition (18) would be violated.
In contrast, the generation of the Uhlmann phase inherently relies on the phase factor $\mathcal{U}$ according to the underlying fiber-bundle structure. All purifications of a given density matrix $\rho$ thus span a U$(N)$ fiber space. Consequently, the Uhlmann phase is naturally and directly connected to the topological properties of the U$(N)$ principal bundle, a feature absent in the IGP. Explicitly, the Uhlmann phase corresponds to the Uhlmann holonomy of the Uhlmann bundle of density matrices, similar to the Berry phase revealing the Berry holonomy of the fiber bundle of pure states Guo et al. (2020). Moreover, the Uhlmann phase requires a cyclic process but is guaranteed to be gauge invariant Sjöqvist et al. (2000); Hou et al. (2023a) while the IGP and its generalization do not impose those constraints. In the following, we will compare the differences between the Uhlmann phase and IGP through specific examples of the CSS and SSS.